OpenStudy (anonymous):

Which conditional and its converse form a true biconditional? A) If x>0, then lxl >0 B) If x=3, then x2^2=9 C) If x^2 =4, Then x=2 D) If x=19, then 2x-3=35

terenzreignz (terenzreignz):

It's basically asking which pair of concepts are essentially saying the same thing... Take choice A for instance... Is it true that if |x| > 0, x > 0? The answer is no, for what if x is negative? Say, -3 |-3| = 3 > 0 But -3 is not greater than 0, so 3 > 0 is false. So it isn't A. Then, which is it? ;)

OpenStudy (anonymous):

C?

terenzreignz (terenzreignz):

So, you're saying, if x^2 is 4, then x = 2?

OpenStudy (anonymous):

yes because 2x2=4

terenzreignz (terenzreignz):

What about -2?

terenzreignz (terenzreignz):

(-2)^2 = 4, but -2 is not 2 :P

OpenStudy (anonymous):

dang then its might be..... B beacuse 3x3x3= 9

terenzreignz (terenzreignz):

First of all, 3x3x3 = 27, not 9 :P

terenzreignz (terenzreignz):

And second, this is similar to choice c But then again, what about -3? (-3)^2 = 9 too, but -3 is not 3...

OpenStudy (anonymous):

i was thinking of addition my bad :x

OpenStudy (anonymous):

wher is the negative comming from?

terenzreignz (terenzreignz):

I mean, just because x^2 = 9 doesn't mean x = 3 Because x could be -3.

OpenStudy (anonymous):

oh then the only answer left is D

terenzreignz (terenzreignz):

Yeah, that's right :P And you just had to try every wrong answer to get it... practice your logic next time, okay? :)

OpenStudy (anonymous):

thanxs ill try thxs for the help